OpenStudy (boldgeorge):

Medal for anyone who can help me with this! An apartment building contains 80 apartments. One bedroom apartments rent for $850 per month. Two bedroom apartments rent for $1050 per month. The total monthly rental income is $78,400. How many apartments of each type are in the building?

OpenStudy (johnweldon1993):

so there are 2 different types let 'x' = 1 bedroom and 'y' = 2 bedroom x + y = 80 1 bedrooms are 850 a month...and 2 bedrooms are 1050 a month 850x + 1050y = 78400 so you have a system of equations x + y = 80 850x + 1050y = 78400 can you solve this?

OpenStudy (texaschic101):

good explaining @johnweldon1993 ...well done

OpenStudy (johnweldon1993):

Thank you! @texaschic101

OpenStudy (texaschic101):

Just telling the truth :)